### v1

@@ 5 Future directions
The development of foundation model technology has achieved remarkable progress, but with the continuous advancement of technology, new challenges and issues arise, pointing out the future research directions.
@@ 5.1 Challenges in data
Acquiring domain-specific data and modeling data structures are the top priority and most important challenges. Foundation models typically require vast amounts of high-quality data for training [116], which may be costly and time-consuming to obtain in specific domains. Moreover, strengthening privacy regulations impose more restrictions on data collection and usage. To address this challenge, future research can focus on developing new data collection and annotation techniques, as well as utilizing synthetic data and weakly supervised learning methods to reduce reliance on extensively labeled data [117, 118]. This not only reduces costs but also effectively utilizes data resources while ensuring privacy protection. On the other hand, effective modeling of data structures in specific domains requires researchers and practitioners to deeply understand domain-specific business processes, extract key business data, and establish comprehensive data preprocessing pipelines.
Understanding multi-modality data poses another critical challenge. Although existing foundation models excel in processing textual data, their understanding of other modalities such as images and sounds still needs improvement [36], not to mention various new data modalities that may emerge in specific domains. Constructing unified models capable of comprehensively processing various modalities of data is essential for enhancing the model’s performance and generalization on multi-modality tasks. This requires researchers and practitioners to not only deeply understand the characteristics of different modalities of data but also explore effective multi-modality fusion and interaction understanding mechanisms.
The availability of aligned multi-modality data presents a potential risk for the long-term development of multi-modal models. While techniques like bridging alignment can reduce the dependency on aligned data [25, 20], the performance and generalization capabilities of multi-modal foundation models still heavily rely on annotated data. This data is costly in terms of both time and financial resources. Therefore, new alignment methods are necessary to address these challenges [53].
@@ 5.2 Challenges in model architecture
In the architectural design of foundation models in specific domains, a core challenge is how to build models that can effectively capture and express deep semantic information specific to different domain [25, 20, 26]. This requires models to not only have a broad knowledge base but also understand and adapt to domain-specific knowledge and input modalities. Foundation models tailored for specific domains need to achieve high modularity and customizability in architecture, enabling adjustment and optimization according to specific application scenarios to adapt to the data characteristics and task requirements of different domains [7, 4, 16]. Furthermore, interpretability of models is particularly important in specific domains. When designing architectures, researchers need to consider how to construct models so that their decision-making processes and output results can be understood and trusted by domain experts and end-users. This may involve developing new model mechanisms, introducing interpretable model intermediate representations, or designing visualization tools to demonstrate the internal workings of the model.
5.3 Challenges in computational resources
Computational resources are also a significant challenge for foundation model technology. Training and running foundation models require enormous computational resources, which not only increases economic costs but also may have environmental impacts. Therefore, researching how to improve the efficiency of model training and inference [119] and how to reduce energy consumption has become an urgent issue. Future research directions may include developing more efficient model compression and acceleration techniques, such as knowledge distillation, model pruning, quantization, etc., as well as exploring more efficient training algorithms and specialized hardware designs.
@@ 5.3 Challenges in computational resources
Lightweight deployment is another important direction for foundation model technology. The size and computational requirements of foundation models often make it difficult to deploy them in mobile devices and edge computational scenarios [120]. To enable foundation models to run in resource-constrained scenarios, lightweight model architectures and deployment strategies need to be developed. This may involve simplifying, distilling, and optimizing models to reduce their size and computational requirements while maintaining or improving their performance. Alternatively, employing cloud-edge collaboration strategies can distribute the training and inference processes of foundation models across various server tiers, enabling a cooperative deployment. In cloud-edge learning, it is essential to consider the joint optimization of sensing, computation, and communication [121, 122, 123]. Such joint optimization for the performance of neural networks and resource consumption holds vast potential for innovation in research and can generate significant value in practical application scenarios.
@@ 5.4 Challenges in security
Security issues are also challenges that foundation model technology must face. Foundation models may be used to generate false information, infringe on privacy, or be maliciously exploited, and the models themselves may also be subject to adversarial attacks [124]. To ensure the security and reliability of models, relevant work includes but is not limited to the following key points: Firstly, strengthening the robustness of models to resist potential adversarial attacks [125]. This may involve developing advanced adversarial training techniques and implementing more stringent data cleaning and preprocessing steps. Secondly, developing and deploying efficient malicious input detection mechanisms, using anomaly detection algorithms and real-time monitoring systems to identify and prevent malicious behavior. Furthermore, ensuring privacy protection is achieved by diminishing the model’s reliance on sensitive data, thereby safeguarding the security and privacy of user information [126]. In addition to technical efforts, enhancing security in workflow and policy aspects is also necessary. For example, implementing security audit and certification processes to comprehensively assess the security of models and ensure that model development and deployment comply with ethical and legal standards, continuously updating security policies to address emerging security threats and challenges.

### v2

@@ 5 Challenges ahead
The development of FM technology has achieved remarkable progress, but with the continuous advancement of technology, new challenges and issues arise, pointing out future research directions.
@@ 5.1 Challenges in data
Acquiring domain-specific data and modeling data structures are the top priority and most important challenges. FMs typically require vast amounts of high-quality data for training [166], which may be costly and time-consuming to obtain in specific domains. Moreover, strengthening privacy regulations imposes more restrictions on data collection and usage. To address this challenge, future research can focus on developing new data collection and annotation techniques and utilizing synthetic data and weakly supervised learning methods to reduce reliance on extensively labeled data [167, 168]. This reduces costs and effectively utilizes data resources while ensuring privacy protection. On the other hand, effective modeling of data structures in specific domains requires researchers and practitioners to deeply understand domain-specific business processes, extract key business data, and establish comprehensive data preprocessing pipelines.
Understanding multi-modality data poses another critical challenge. Although existing FMs excel in processing textual data, their understanding of other modalities, such as images and sounds, still needs improvement [36], not to mention various new data modalities that may emerge in specific domains. Constructing unified models capable of comprehensively processing various modalities of data is essential for enhancing the model’s performance and generalization on multi-modality tasks. This requires researchers and practitioners to not only deeply understand the characteristics of different modalities of data but also explore effective multi-modality fusion and interaction understanding mechanisms.
The availability of aligned multi-modality data presents a potential risk for the long-term development of multi-modal models. While techniques like bridging alignment can reduce the dependency on aligned data [25, 20], the performance and generalization capabilities of multi-modal FMs still heavily rely on annotated data. This data is costly in terms of both time and financial resources. Therefore, new alignment methods are necessary to address these challenges [52].
@@ 5.2 Challenges in model architecture
In the architectural design of FMs in specific domains, a core challenge is how to build models that can effectively capture and express deep semantic information specific to different domain [25, 20, 26]. This requires models to have a broad knowledge base and understand and adapt to domain-specific knowledge and input modalities. FMs tailored for specific domains need to achieve high modularity and customizability in architecture, enabling adjustment and optimization according to specific application scenarios to adapt to the data characteristics and task requirements of different domains [7, 4, 16]. Furthermore, interpretability of models is particularly important in specific domains. When designing architectures, researchers need to consider how to construct models so that their decision-making processes and output results can be understood and trusted by domain experts and end-users. This may involve developing new model mechanisms, introducing interpretable model intermediate representations, or designing visualization tools to demonstrate the internal workings of the model.
5.3 Challenges in model training and deployment cost
The demands on computational, memory, and energy resources put significant limitations on the development of FM technologies. Therefore, their cost-intensive nature should be primarily considered when designing domain-specific FMs. Research on improving the efficiency of model training and inference [169] and reducing resource consumption has become an urgent issue. Potential research directions include developing more efficient memory-saving and acceleration techniques, such as quantization, efficient attention, distributed learning, etc. These technologies can help model deployers reduce costs when maintaining model performance while building and deploying a domain-specific FM.
5.3.1 Challenges in cost-efficient training
The impressive capabilities of FMs come at a significant cost of computing, memory, and energy during training. For example, in some Internet of Things (IoT) scenarios, such costs could be the primary concern for model deployers. On one hand, they may need to use edge devices to collect real-time data for training large models rather than centrally pre-training a model and then deploying it. On the other hand, the resource demands of training models pose a hard constraint on edge devices. Since building domain-specific FMs is costly, it is beneficial to help model deployers better manage their costs.
Memory efficiency: FMs’ huge parameter scale put a serious challenge for memory resources. Mixed precision training [170] is a common technique to convert parts of high-precision variables into low-precision dynamically implemented by automatic mixed precision [171] in each training step. This can accelerate model training and reduce GPU memory usage while minimally impacting model accuracy. A huge memory is occupied while focusing on masked training in long-sequence training. Therefore, positional encoding compression methods like ALiBi [172] and reset attention mask [173] have been proposed.
Computation efficiency: Most existing FMs have transformer-based architecture in which the attention mechanism has a powerful capability but is also computationally intensive. Researchers are now working to create efficient transformers. FlashAttention [174] significantly improves the efficiency of attention computation by IO-awareness block attention algorithm, and researchers also proposed other attention algorithms like LSH attention [175] and local attention [111]. Parallel training is another computational accelerating approach, including traditional data parallel and emerging methods like ZeRO method [176] distribute data and model parameters to achieve fast training under limited hardware constraints. Additionally, parameter-efficient fine-tuning methods [103, 107] are also the key approach to improve the computation efficiency.
Communication efficiency: Federated learning [177] is applied when the training data cannot be acquired and centralized in advance due to privacy concerns. Reducing communication overhead is essential since FL involves frequent model parameter exchanges. Parameter-efficient fine-tuning methods have been combined with federated learning for efficient communication [178, 179, 180, 181].
Energy efficiency: The energy consumption of FM training is an emerging area as it generates significant carbon footprints [182], leading to environmental impacts at the same time as increasing economic costs.
5.3.2 Challenges in cost-efficient deployment
The deployment of domain-specific FMs requires enormous resources, not only the training process but also the deployment of domain-specific FMs. In domain-specific business scenarios, the user’s requirements for model response speed and performance may be an important consideration. To meet these demands under the resource constraints on the hardware, cost-efficient deployment strategies are highly desired, as introduced below:
Inference efficiency: The inference time of domain-specific FMs drastically increases with a huge amount of data. To solve such a problem, inference acceleration is crucial for deploying domain-specific FMs. KV cache [183] is a popular method to speed up FM inference by only generating a single row of the Q matrix. Furthermore, multi-query attention [184] could reduce the size of the KV cache, and paged attention [185] utilizes the block table to make full use of the large KV cache. In addition to the improvement on transformer architecture, speculative decoding [186] generates the next token with the help of a small “draft” model, which provides a set of token candidates for the main FM, followed by various methods such as staged speculative decoding [187], guided generation [188], lookahead decoding [189] and prompt lookup decoding [190]. Additionally, edge AI draws researchers’ attention as it reduces the latency of providing AI services to devices [191, 192, 193].
@@ 5.3.2 Challenges in cost-efficient deployment
Resource efficiency: The resource requirements of FMs often make it difficult to deploy them in mobile devices and edge computational scenarios [194]. To enable FMs to run in resource-constrained scenarios, lightweight model architectures and deployment strategies need to be developed. This may involve simplifying, distilling, and optimizing models to reduce their resource requirements while maintaining or improving their performance. Alternatively, employing cloud-edge collaboration strategies can distribute the training and inference processes of FMs across various server tiers, enabling a cooperative deployment. In cloud-edge learning, it is essential to consider the joint optimization of sensing, computation, and communication [195, 196, 197]. Such joint optimization for the performance of neural networks and resource consumption holds vast potential for innovation in research and can generate significant value in practical application scenarios. Furthermore, energy consumption in the inference phase is becoming the new trend within the community [198]. There are trade-offs between performance and power usage to achieve energy-efficient FM inference [199, 200, 201, 202]. Recently, edge AI inference [203, 204] also shows potential to enable various devices without adequate computing resources to access the FMs.
@@ 5.4 Challenges in security
Security issues are also challenges that FM technology must face. FMs may be used to generate false information, infringe on privacy, or be maliciously exploited, and the models themselves may also be subject to adversarial attacks [205]. To ensure the security and reliability of models, relevant work includes but is not limited to the following key points: Firstly, strengthening the robustness of models to resist potential adversarial attacks [206]. This may involve developing advanced adversarial training techniques and implementing more stringent data cleaning and preprocessing steps. In addition to traditional adversarial training, a representative line of early work proposed certified robustness methods—such as randomized smoothing [207]—which provide guarantees against perturbations. More recent studies have explored robust fine-tuning strategies, including adversarial contrastive learning and consistency regularization, to enhance generalization under distribution shifts. In parallel, robust representation learning has emerged as a promising direction, aiming to learn perturbation-invariant and semantically consistent embeddings, as exemplified by adversarial contrastive objectives in robust pre-training [208]. These complementary techniques contribute to a layered defense system that strengthens the reliability of foundation models in open environments. Secondly, developing and deploying efficient malicious input detection mechanisms, using anomaly detection algorithms and real-time monitoring systems to identify and prevent malicious behavior. Furthermore, ensuring privacy protection is achieved by diminishing the model’s reliance on sensitive data, thereby safeguarding the security and privacy of user information [209]. In addition to technical efforts, enhancing security in workflow and policy aspects is also necessary. For example, implementing security audit and certification processes to comprehensively assess the security of models and ensure that model development and deployment comply with ethical and legal standards, continuously updating security policies to address emerging security threats and challenges.
6 Conclusion
This paper provides a comprehensive overview of recent advancements in domain-specific foundation models (FMs). We begin by discussing the foundational concepts of FMs, including their architectures, training methodologies, the benefits of scaling, and a comparison of their performance across different contexts. Next, we delve into the key technologies for developing domain-specific FMs, which encompass three primary approaches: enhancing general-purpose FMs with domain-specific knowledge, customizing FMs using pre-trained modules, and building FMs from scratch without relying on pre-trained components. We then explore the diverse applications of domain-specific FMs across various fields, including telecommunications, autonomous driving, mathematics, medicine, law, the arts, and finance. Following this, we address the significant challenges faced by practitioners in constructing and deploying domain-specific FMs. These challenges span multiple dimensions, such as data availability and quality, model architecture design, training processes, deployment complexities, and security concerns. By synthesizing these insights, we aim to provide researchers and practitioners with a valuable resource for navigating the development and application of domain-specific FMs. We hope this paper serves as a foundation to guide future innovation and encourage the creation of tailored FMs that address specific domain requirements effectively.
[…]
This work was supported in part by National Natural Science Foundation of China (Grant No. 62371313), supported by the National Science and Technology Major Project (No. 2024ZD1300500), in part by Guangdong Major Project of Basic and Applied Basic Research (No. 2023B0303000001), in part by Guangdong Young Talent Research Project (Grant No. 2023TQ07A708), in part by ShenzhenHong Kong-Macau Technology Research Programme (Type C) (Grant No. SGDX202308 21091559018), in part by the Shenzhen Science and Technology Program (Grant No. JCYJ20241202124934046), funded by Young Elite Scientists Sponsorship Program by CAST (2022QNRC001).